OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2 y- 9 = 2(x -5) y -5 = 2(x + 9) y + 9 = 2(x -5) y -9 = 2(x + 5)

OpenStudy (anonymous):

y-9=2(x-5)

OpenStudy (anonymous):

as equation for st. line is slope-point form is as - y-y1=m(x-x1) where x1,y1 are given points ....................